I am desperately trying to expain how to predict and test problem solving,is only a third grader and has not been taught algebra so not able to explain that way. problems are as such... Peter delievers 110 newspapers on the weekend. He delivers 20 more papers on Sunday than on Saturday. How many papers does he deliver each day?
Answer by stanbon(26297)

Peter delivers 110 newspapers on the weekend. He delivers 20 more papers on Sunday than on Saturday. How many papers does he deliver each day?
-----------------------------
Let number delivered on Saturday be "x".
Then number delivered on Sunday is "x+20".
-----------------
Equation:
x + x+20 = 110
2x + 20 = 110
2x = 90
x = 45 (# he delivers on Saturday)
x+20 = 65 (# he delivers on Sunday)
======================================
Cheers,
Stan H.